OpenStudy (anonymous):

2q + 18 = –5q – 3 A. q = –5 B. q = –3 C. q = 3 D. q = 5

OpenStudy (kendricklamar2014):

B. Q = -3

OpenStudy (anonymous):

how?

OpenStudy (texaschic101):

2q + 18 = -5q - 3 -- subtract 18 2q = -5q - 3 - 18 --- add 5q 2q + 5q = -3 - 18 -- combine like terms 7q = -21 -- divide by 7 q = -3